OpenStudy (anonymous):

-6p - 21 = 3p - 12 A. 1 B. 3 C. –3 D. –1

OpenStudy (anonymous):

-6p - 21 = 3p - 12 -3 p -3p -9p -21 = -12 +21 +21 -9p = 9 p=-1